# Advancing Brainwave-Based Biometrics: A Large-Scale, Multi-Session Evaluation

### v2

@@ I Introduction
  • Channel Reduction (Sec IV-D): To evaluate the system’s performance under conditions resembling real-world scenarios, we reduced the number of channels in the PEERS dataset—collected using medical-grade devices—from 93 to 14, 7, and 4. These reductions correspond to sensor placements commonly found in consumer-grade EEG devices. The results showed a relatively small increase in error compared to the notable reduction in channels.
@@ III-B Preprocessing
As the channel layout of the different headsets varied, we unified them by mapping the channels to a well-known extended version [^4] https://mne.tools/stable/auto_tutorials/intro/40_sensor_locations.html of the ten-twenty electrode system [31], which includes 93 channels. Subsequently, the EEG data channels were reordered for each recording to align with the expected order during evaluation across all three devices. Following this, and in line with best practices for preprocessing in brainwave authentication [14, 5, 32, 12], a common average reference (CAR) was applied to reduce spatially correlated noise and improve signal quality. Next, we applied a bandpass filter between 1.0 Hz and 50.0 Hz using a zero-phase finite impulse response (FIR) filter with default parameters, as implemented in MNE-Python [33], to focus on frequency bands that contain EEG data relevant for authentication [34]. As the 50 Hz power-line frequency lies close to the upper end of the preserved frequency band, we additionally applied a notch filter at 50 Hz to suppress potential residual line noise. After filtering, the data was normalized using a robust normalization approach, where the median was subtracted and the result divided by the interquartile range (IQR), mitigating the influence of outliers while preserving the temporal structure of the signal. The data then consisted of 1-second non-overlapping samples aligned to stimulus timestamps. Each sample was stored together with metadata including session date and hardware type.
@@ III-C Feature Extraction
For each loss function, different encoder architectures were explored to optimize feature extraction. Specifically, we implemented six backbone EEG models — ResNet1D [43], ShallowNet [44], DeepConvNet [44], EEGNet [45], LSTM [46], and GRU [46] — as encoder networks. These models were selected to cover a diverse range of architectures, including convolutional and recurrent networks, and to balance model complexity and representational capacity for EEG signal processing.
@@ III-F Performance Metrics
TABLE II: Results of top 1,000 hyperparameter trials on validation set (15 subjects): pipeline counts, average EER, and best EER among all trials.
@@
Feature Extraction
ResNet1D + SupConLoss 98 10.20 8.05
@@
DeepConvNet + SoftTripletLoss 90 12.19 10.23
GRU + SoftTripletLoss 36 11.43 10.35
LSTM + SoftTripletLoss 28 11.87 10.72
@@
EEGNet + ArcFaceLoss 0 0.00 13.18
LSTM + ArcFaceLoss 0 0.00 13.21
EEGNet + SupConLoss 0 0.00 13.36
LSTM + SupConLoss 0 0.00 13.61
DeepConvNet + TripletMarginLoss 0 0.00 13.88
EEGNet + SoftTripletLoss 0 0.00 14.49
EEGNet + TripletMarginLoss 0 0.00 15.00
LSTM + TripletMarginLoss 0 0.00 15.55
GRU + TripletMarginLoss 0 0.00 18.74
PSD + AR 0 0.00 19.23
AR 0 0.00 19.43
PSD 0 0.00 29.34
Feature Comparison
Euclidean Distance 86 11.18 8.05
@@ IV-A Benchmarking Feature Extraction and Compression
Experiment setup: Our claim is that the proposed method supports both inter-subject and inter-session generalization. Therefore, this is reflected by the data split and evaluation protocol. We divided the 345 available subjects into non-overlapping sets: 230 subjects for training the deep learning-based feature extractor, 15 for validation, and 100 randomly selected subjects for evaluation, on whom system performance is measured. In all experiments, unless specified otherwise, the first session of each subject was used for enrollment, and the remaining sessions of the same subject for verification. This setup ensures inter-subject evaluation and shows that the model does not require retraining when new users are added. Additionally, separating enrollment and verification by session enables the assessment of inter-session generalization.
Hyperparameter tuning: From each combination of model architectures and loss functions (6×5) mentioned in the methods (Section III), we trained a model using hyperparameter tuning on the optimizer, learning rate, batch size, and embedding dimension based on the training set, guided by performance on the validation set. The entire process took approximately 144 hours on four NVIDIA GeForce RTX 4090 GPUs.
@@
Feature Extraction Methods: Table II shows that our dataset is sufficient for deep learning approaches to outperform handcrafted features, which result in significantly higher error. Among the encoders, ResNet1D consistently achieved the top three ranks, followed by ShallowNet, which outperformed the remaining models. This suggests that convolutional networks generally perform better than recurrent architectures. Among the convolutional models, simpler architectures like ShallowNet may benefit from reduced overfitting, while ResNet1D leverages residual connections to achieve better performance than models such as DeepConvNet. Although EEGNet is relatively shallow, its specialized filters for frequency and spatial information did not translate into strong performance in our evaluation. Regarding loss functions, although SoftTripleLoss—with its multiple learnable centers per class—performed best with shallower and recurrent backbones (ShallowNet, DeepConvNet, LSTM, GRU), ResNet1D combined with SupConLoss achieved the highest overall performance across all architectures. SupConLoss combines the strengths of distance-based and softmax-based approaches by leveraging all positive and negative pairs in the batch, guided by class labels. This batch-level supervision enforces strong intra-class cohesion and inter-class separation, which ResNet1D—with its higher capacity—can fully exploit to learn highly discriminative EEG embeddings
Comparison Methods: The results (Table II) suggest that, in general, distance-based functions outperform shallow classifier approaches for feature comparison. This is not surprising, as loss functions in metric learning are designed to minimize distances between embeddings, typically using cosine or Euclidean distance. In contrast, shallow classifiers either require negative samples for training or, in the case of one-class classification, often show relatively poor performance due to the instability associated with one-class models. This is also why state-of-the-art face recognition methods use distance functions for feature comparison.
@@
Figure 2: EER vs. Number of Subjects: Each blue dot represents the average of 50 instances of randomly selecting $N$ subjects and calculating the EER. The dot size represents the standard deviation of the EER values across these 50 calculations.
@@ IV-B3 Visualize Embedding Space
To investigate the embedding space while accounting for color and spatial limitations, we randomly selected 20 subjects to prevent overcrowding the embedding space with excessive data. [^7] BioSemi subjects were excluded to avoid large gaps in the plot and enhance detail visibility. Figure 6 illustrates the embedding space for these selected subjects, demonstrating that the feature extraction method effectively identifies subject identities. However, a session effect is evident, where samples from the same session tend to form distinct clusters. To better visualize this phenomenon, the upper-right region of Figure 6 is magnified in Figure 7, with separate colors indicating different sessions, further confirming the session-based clustering. The session effect on clustering may arise from environmental noise, changes in hair length, slight variations in the positioning of brainwave channels on the head, hardware differences, and natural changes in brainwave patterns over time due to factors such as aging, personal experiences, or mental states (e.g., stress or focus levels) throughout sessions.
@@ IV-B4 Impact of Different Headsets on Performance
Cross Device: In real-world scenarios, individuals using brainwave-based authentication systems may rely on different recording devices. Therefore, authentication systems should maintain performance when enrollment and verification data come from different hardware. To evaluate this, we filtered genuine scores based on unique enrollment-verification device pairs and calculated the EER for each cross-device setting. The results show an EER of 12.53% for Geodesic–HydroCe pairs and 12.56% for HydroCe–Geodesic pairs. In contrast, BioSemi–HydroCe and HydroCe–BioSemi pairs yield much higher EERs of 39.00% and 44.05%, respectively. For other combinations (e.g., Geodesic–BioSemi), no data were available in the evaluation set. The reason for this variation is that in the training set used for feature extraction, only 7 subjects had HydroCe–BioSemi session pairs and only 2 had BioSemi–HydroCe pairs out of 230 subjects, while 122 subjects had Geodesic–HydroCe session pairs. This limited overlap likely contributes to the poor generalization in the BioSemi–HydroCe condition and explains the formation of separate clusters in the t-SNE visualization shown in Figure 5, where BioSemi forms a distinct cluster.
@@
[…]
@@
[…]
@@ IV-C Enhancing Performance
Samples from the same session are relatively close in the embedding space, but in some cases, they are more dispersed. For example, the pink session in Figure 6. Therefore utilizing multiple verification samples can improve results, albeit at the cost of additional verification time. Table VII illustrates the tradeoff between the number of verification samples, EER, and verification time. The results show that using more samples reduces the error, but improvements beyond four samples show a slower pace in reducing error. Considering the time cost of approximately one second per sample, using four samples appears to be a reasonable balance for the averaging strategy.
TABLE VI: Summary of state-of-the-art brain biometric verification solutions evaluated on multisession datasets
@@
[…]
Proposed Method 345 6007 (17$\sim$ 17∼ 17) up to 66 months 1s 10.28
Moreover, while comparing the enrollment session to all verification sessions offers useful insights into brainwave authentication, real-world scenarios typically involve a single initial enrollment session followed by one verification attempt after some time. Upon successful login, it may be possible to update the enrollment set with new data from the verification attempt, which can then be used in future verifications. Therefore, in Table VIII, we consider the first session as enrollment and the second as verification and report the results. We observe an EER of 6.87% based on single verification sample and 3.04% for four verification samples, showing significant improvement compared to the previous 10.28% and 6.22%. Additionally, we investigate the effect of multiple enrollment sessions. The results suggest that having two enrollment sessions improves EER to 4.87% and 1.54%, but further enrollment sessions show diminishing returns. The average time interval between the two enrollment sessions was 28 days. Among the subjects, 88 had a session interval of 10 days or less, and 32 had intervals of only one or two days.
TABLE VII: Tradeoff between the number of samples (SN) and EER. Increasing the number of samples reduces the EER but increases the verification time, as shown in the table.
@@
TABLE VIII: Tradeoff between the number of enrollment sessions (ES) and EER when using 1 verification sample (1V) and 4 verification samples (4V). The table also includes the number of available subjects (NAS) used in the evaluation.
@@
TABLE IX: EER (%) when using the same channels as consumer-grade devices. $E_{x}$ denotes the number of enrollment sessions, and $xV$ denotes the number of verification samples. For the first two columns (Baseline and 4V), all possible verification sessions were considered. However, for the remaining columns, only the first verification session immediately following the last enrollment session was used.
@@
IV-D Practicality with Consumer-Grade Channels

### v3

@@ I Introduction
  • Channel Reduction (Sec X): To evaluate the system’s performance under conditions resembling real-world scenarios, we reduced the number of channels in the PEERS dataset—collected using medical-grade devices—from 93 to 14, 7, and 4. These reductions correspond to sensor placements commonly found in consumer-grade EEG devices. The results showed a relatively small increase in error compared to the notable reduction in channels.
@@ III-B Preprocessing
As the channel layout of the different headsets varied, we unified them by mapping the channels to a well-known extended version [^4] https://mne.tools/stable/auto_tutorials/intro/40_sensor_locations.html of the ten-twenty electrode system [53], which includes 93 channels. Subsequently, the EEG data channels were reordered for each recording to align with the expected order during evaluation across all three devices. Following this, and in line with best practices for preprocessing in brainwave authentication [6, 3, 45, 35], a common average reference (CAR) was applied to reduce spatially correlated noise and improve signal quality. Next, we applied a bandpass filter between 1.0 Hz and 50.0 Hz using a zero-phase finite impulse response (FIR) filter with default parameters, as implemented in MNE-Python [14], to focus on frequency bands that contain EEG data relevant for authentication [55]. As the 50 Hz power-line frequency lies close to the upper end of the preserved frequency band, we additionally applied a notch filter at 50 Hz to suppress potential residual line noise  [^5] Residual power-line interference in EEG recordings typically occurs at 50 Hz (e.g., most of Europe) or 60 Hz (e.g., North America). We bandpass-filtered the data between 1 and 50 Hz and then applied notch filters at 50 Hz and 60 Hz to suppress residual line noise at both mains frequencies; because 60 Hz lies outside the preserved band, its notch filter effect on the data is negligible.. After filtering, the data was normalized using a robust normalization approach, where the median was subtracted and the result divided by the interquartile range (IQR), mitigating the influence of outliers while preserving the temporal structure of the signal. The data then consisted of 1-second non-overlapping samples aligned to stimulus timestamps. Each sample was stored together with metadata including session date and hardware type.
@@ III-C Feature Extraction
For Triplet Loss, we adopt semi-hard online triplet mining [51], which is recommended to emphasize informative (non-trivial) triplets and improve convergence. LiftedStructureLoss and SupConLoss are pair-based (and not triplet-based); accordingly, we employ Multi-Similarity mining [57] to select hard positive and negative pairs based on cosine similarity. ArcFaceLoss and SoftTripleLoss are proxy-based margin losses (softmax-based) that operate on class centers and therefore do not require mining, as they do not consume explicit pair or triplet indices. Samples are drawn across all sessions in the training set, and positive pairs include both within-session and cross-session pairs of the same subject. This allows the model to minimize both intra-session and inter-session embedding distances, supporting continuous authentication (stability within sessions) as well as multi-session authentication (robustness across sessions).
For each loss function, different encoder architectures were explored to optimize feature extraction. Specifically, we implemented six backbone EEG models — ResNet1D [62], ShallowNet [49], DeepConvNet [49], EEGNet [32], LSTM [61], and GRU [61] — as encoder networks. These models were selected to cover a diverse range of architectures, including convolutional and recurrent networks, and to balance model complexity and representational capacity for EEG signal processing. Details of the neural network architecture are in Appendix B.
@@ III-F Performance Metrics
TABLE II: Results of hyperparameter trials on the validation set (15 subjects): number of pipeline occurrences within the top 1,000 trials, average EER computed for pipelines appearing in the top 1,000, and best EER across all 10,000 trials.
@@
Feature Extraction ResNet1D + SupConLoss 98 10.20 8.05
@@
DeepConvNet + SoftTripleLoss 90 12.19 10.23
GRU + SoftTripleLoss 36 11.43 10.35
LSTM + SoftTripleLoss 28 11.87 10.72
@@
EEGNet + ArcFaceLoss 0 13.18
LSTM + ArcFaceLoss 0 13.21
EEGNet + SupConLoss 0 13.36
LSTM + SupConLoss 0 13.61
DeepConvNet + TripletMarginLoss 0 13.88
EEGNet + SoftTripleLoss 0 14.49
EEGNet + TripletMarginLoss 0 15.00
LSTM + TripletMarginLoss 0 15.55
GRU + TripletMarginLoss 0 18.74
PSD + AR 0 19.23
AR 0 19.43
PSD 0 29.34
Feature Comparison Euclidean Distance 86 11.18 8.05
@@ IV-A Benchmarking Feature Extraction and Compression
IV-A1 Experiment setup
We hypothesize that the proposed method supports both inter-subject and inter-session generalization. Inter-subject generalization refers to the ability of a model trained on a subset of users to perform well on previously unseen users. Inter-session generalization refers to the ability of the model to embed data from different sessions of the same user close to each other in the feature space. Therefore, this is reflected by the data split and evaluation protocol. We divided the 345 available subjects into non-overlapping sets: 230 subjects for training the deep learning-based feature extractor, 15 for validation, and 100 randomly selected subjects for evaluation, on whom system performance is measured. In all experiments, unless specified otherwise, the first session of each subject in the evaluation set was used for enrollment, and the remaining sessions of the same subject for verification. This split ensures inter-subject evaluation, since no subject appears in more than one of the training, validation, and evaluation sets, and shows that the model does not require retraining when new users are added. Additionally, separating enrollment and verification by session enables the assessment of inter-session generalization.
IV-A2 Hyperparameter tuning
From each combination of model architectures and loss functions (6×5) mentioned in the methods (Section III), we trained a model using hyperparameter tuning on the optimizer, learning rate, batch size, and embedding dimension based on the training set, guided by performance on the validation set. The entire process took approximately 144 hours on four NVIDIA GeForce RTX 4090 GPUs.
@@ IV-A2 Hyperparameter tuning
IV-A3 Feature Extraction Methods
Table II shows that our dataset is sufficient for deep learning approaches to outperform handcrafted features, which result in significantly higher error. Among the encoders, ResNet1D consistently achieved the top three ranks, followed by ShallowNet, which outperformed the remaining models. This suggests that convolutional networks generally perform better than recurrent architectures. Among the convolutional models, simpler architectures like ShallowNet may benefit from reduced overfitting, while ResNet1D leverages residual connections to achieve better performance than models such as DeepConvNet. Although EEGNet is relatively shallow, its specialized filters for frequency and spatial information did not translate into strong performance in our evaluation. Regarding loss functions, although SoftTripleLoss—with its multiple learnable centers per class—performed best with shallower and recurrent backbones (ShallowNet, DeepConvNet, LSTM, GRU), ResNet1D combined with SupConLoss achieved the highest overall performance across all architectures. SupConLoss combines the strengths of distance-based and softmax-based approaches by leveraging all positive and negative pairs in the batch, guided by class labels. This batch-level supervision enforces strong intra-class cohesion and inter-class separation, which ResNet1D—with its higher capacity—can fully exploit to learn highly discriminative EEG embeddings
IV-A4 Comparison Methods
The results (Table II) suggest that, in general, distance-based functions outperform shallow classifier approaches for feature comparison. This is not surprising, as loss functions in metric learning are designed to minimize distances between embeddings, typically using cosine or Euclidean distance. In contrast, shallow classifiers either require negative samples for training or, in the case of one-class classification, often show relatively poor performance due to the instability associated with one-class models. This is also why state-of-the-art face recognition methods use distance functions for feature comparison.
@@ IV-A4 Comparison Methods
Figure 2: EER vs. Number of Subjects: Each blue dot represents the average of 50 instances of randomly selecting NN subjects and calculating the EER. The dot size represents the standard deviation of the EER values across these 50 calculations.
@@ IV-B3 Visualize Embedding Space
To investigate the embedding space while accounting for color and spatial limitations, we randomly selected 20 subjects to prevent overcrowding the embedding space with excessive data. [^8] BioSemi subjects were excluded to avoid large gaps in the plot and enhance detail visibility. Figure 6 illustrates the embedding space for these selected subjects, demonstrating that the feature extraction method effectively identifies subject identities. However, a session effect is evident, where samples from the same session tend to form distinct clusters. To better visualize this phenomenon, the upper-right region of Figure 6 is magnified in Figure 7, with separate colors indicating different sessions, further confirming the session-based clustering. The session effect on clustering may arise from environmental noise, changes in hair length, slight variations in the positioning of brainwave channels on the head, hardware differences, and natural changes in brainwave patterns over time due to factors such as aging, personal experiences, or mental states (e.g., stress or focus levels) throughout sessions. As a consequence, most samples within a session tend to be either close to or far from the enrollment set at the same time, which can contribute to high variability in performance, as reflected in the relatively large standard deviation ($\pm 6.3\%$) observed in the reported EER (10.28%).
@@ IV-B4 Impact of Different Headsets on Performance
 [^9] To clarify, the results in IV.A (Table II) and Sections IV.B.1–IV.B.3 are computed on the combined data from all the three headsets.
@@
Cross Device: In real-world scenarios, individuals using brainwave-based authentication systems may rely on different recording devices. Therefore, authentication systems should maintain performance when enrollment and verification data come from different hardware. To evaluate this, we used the general model trained on the combined training set from all three headsets (230 subjects) and then filtered genuine scores in the evaluation set (100 subjects) by enrollment–verification device pairs to compute the EER for each cross-device setting. The results show an EER of 12.53% for Geodesic–HydroCe pairs and 12.56% for HydroCe–Geodesic pairs. In contrast, BioSemi–HydroCe and HydroCe–BioSemi pairs yield much higher EERs of 39.00% and 44.05%, respectively. For other combinations (e.g., Geodesic–BioSemi), no data were available in the evaluation set. The reason for this variation is that in the training set used for feature extraction, only 7 subjects had HydroCe–BioSemi session pairs and only 2 had BioSemi–HydroCe pairs out of 230 subjects, while 122 subjects had Geodesic–HydroCe session pairs. This limited overlap likely contributes to the poor generalization in the BioSemi–HydroCe condition and explains the formation of separate clusters in the t-SNE visualization shown in Figure 5, where BioSemi forms a distinct cluster.
@@
Geodisi $9.52\pm\text{\scriptsize 6.1}$ $16.06\pm\text{\scriptsize 8.4}$ $29.33\pm\text{\scriptsize 10.9}$ $8.32\pm\text{\scriptsize 5.2}$
HydroCe $18.56\pm\text{\scriptsize 10.0}$ $9.24\pm\text{\scriptsize 7.4}$ $28.32\pm\text{\scriptsize 11.7}$ $9.90\pm\text{\scriptsize 8.4}$
BioSemi $2.63\pm\text{\scriptsize 9.8} 33$$3.35\pm\text{\scriptsize 10.5}$ $23.52\pm\text{\scriptsize 8.4}$ $20.62\pm\text{\scriptsize 7.6}$
All $2.60\pm\text{\scriptsize 12.6} 22$$2.08\pm\text{\scriptsize 14.9}$ $28.70\pm\text{\scriptsize 11.6}$ $10.28\pm\text{\scriptsize 6.3}$
@@
Geodisi $17.37\pm\text{\scriptsize 7.8}$ $16.63\pm\text{\scriptsize 8.6}$ $30.10\pm\text{\scriptsize 11.5}$ $10.55\pm\text{\scriptsize 6.3}$
HydroCe $24.31\pm\text{\scriptsize 9.6}$ $10.91\pm\text{\scriptsize 8.7}$ $29.65\pm\text{\scriptsize 12.2}$ $10.09\pm\text{\scriptsize 8.5}$
BioSemi $33.52\pm\text{\scriptsize 10.0}$ $36.31\pm\text{\scriptsize 11.2}$ $22.71\pm\text{\scriptsize 8.0}$ $21.55\pm\text{\scriptsize 7.9}$
All $24.46\pm\text{\scriptsize 11.5}$ $23.27\pm\text{\scriptsize 14.9}$ $28.22\pm\text{\scriptsize 13.4}$ $21.09\pm\text{\scriptsize 14.4}$
@@ IV-C Enhancing Performance
Samples from the same session are relatively close in the embedding space, but in some cases, they are more dispersed. For example, the pink session in Figure 6. Therefore utilizing multiple verification samples can improve results, albeit at the cost of additional verification time. Table VI illustrates the tradeoff between the number of verification samples, EER, and verification time. The results show that using more samples reduces the error, but improvements beyond four samples show a slower pace in reducing error. Considering the time cost of approximately one second per sample, using four samples appears to be a reasonable balance for the averaging strategy.
@@
TABLE VI: Tradeoff between the number of samples (SN) and EER. Increasing the number of samples reduces the EER but increases the verification time, as shown in the table.
@@
TABLE VII: Tradeoff between the number of enrollment sessions (ES) and EER when using 1 verification sample (1V) and 4 verification samples (4V). The table also includes the number of available subjects (NAS) used in the evaluation.
@@
TABLE VIII: EER (%) when using the same channels as consumer-grade devices. $E_{x}$ denotes the number of enrollment sessions, and xVxV denotes the number of verification samples. For the first two columns (Baseline and 4V), all possible verification sessions were considered. However, for the remaining columns, only the first verification session immediately following the last enrollment session was used.
@@
TABLE IX: Verification performance (EER [%]) for different scalp regions (frontal, central, temporal, parietal–occipital)
Region Baseline 4V E1 1V E1 4V E2 1V E2 4V
Frontal 15.74 10.36 12.51 6.49 10.63 5.13
Central 16.36 11.31 12.41 6.59 10.67 4.94
Temporal 15.76 10.90 11.72 6.48 10.11 4.53
Par.–Occ. 15.11 9.95 11.67 5.82 9.52 4.33
TABLE X: Summary of state-of-the-art brain biometric verification solutions evaluated on multisession datasets
Das et al.  [7], 2016 50 150 (3) up to 49 days 30s $\times$ $\times$ $\times$ $10<\text{EER}<22$
Wu et al.  [59], 2018 40 80 (2) 30 days 6s $\times$ $\times$ $\times$ $5.75<\text{EER}<7.07$
Maiorana et al.  [34], 2018 45 270 (6) up to 36 months 5s $\times$ $\times$ $\times$ $6.6\leq\text{EER}\leq 10.7$
Maiorana [35], 2021 45 225 (5) up to 15 months 5s $\times$ $\times$ $4.8\leq\text{EER}\leq 10.7$
Debie et al. [8], 2021 54 108 (2) different days [^3] The paper did not specify the exact number of days, so we used the same terminology. However, it typically refers to a very short time interval. $7.3s$ $\times$ $\times$ $1.16<\text{EER}<1.93$
Proposed Method 345 6007 ($\sim 17$) up to 66 months 1s 10.28
IV-D Channels Reduction
In this section, we analyze how reducing the number and placement of EEG channels affects biometric authentication performance. We consider two practically relevant scenarios: limiting channels to those available on specific consumer-grade headsets and restricting channels to subsets from distinct scalp regions under a fixed channel budget.
IV-D1 Consumer-grade headsets